OpenStudy (chycora):

Let x represent the measure of angle ACD, and let y represent the measure of angle BCF. The value of x is ° and the value of y is °.

OpenStudy (chycora):

@iGreen. @AnswerMyQuestions

OpenStudy (chycora):

@Austin316

OpenStudy (anonymous):

using the anglesum property u can find angle acb and then subtracting it from 180 will give u x.

OpenStudy (anonymous):

after that dce and acb are vertically opposite and then u can find y by, y+ angle dce+180=0

OpenStudy (chycora):

so 180 - 80=100 180 + 34=214 am i correct

OpenStudy (anonymous):

acb = 180-(90+34) and x = 180- acb.

OpenStudy (anonymous):

dce=acb and then again y + dce + 80 = 180

OpenStudy (chycora):

ok so acb = 180-(90+34) 180- 124 =56

OpenStudy (chycora):

@Godlovesme

OpenStudy (chycora):

ok dce=acb and then again y + dce + 80 = 180 and i don't get this part @Austin316

OpenStudy (godlovesme):

sorry idk this :( @chycora

OpenStudy (chycora):

ur kool

OpenStudy (anonymous):

its beacuse they are forming the linear pair.

OpenStudy (anonymous):

u can see in the figure yourself.

OpenStudy (chycora):

k thanx

OpenStudy (chycora):

i gave u a medal by the way

OpenStudy (anonymous):

thaxxxx for that :)

OpenStudy (chycora):

welcome ur the best lol